OpenStudy (anonymous):

Two trains each make a journey of 330km. One of the train travels 5km/h faster than the other and takes 30 min less time. Find the speed of the trains.

OpenStudy (anonymous):

330 = vt 330 = (v+5)(t - 0.5)

OpenStudy (anonymous):

You have two velocities, \(x\) and \(y\). Our equations are:\[ y=x+5 \]And \[ \frac{x}{330} + \frac{30}{60}= \frac{y}{330} \]

OpenStudy (anonymous):

Whoops, I should have put \[ \frac{330}{x}+\frac{30}{60}=\frac{330}{y} \]Since \(d=vt \implies t=d/v\)

OpenStudy (anonymous):

let speed of first train be s1,and time t1 330=s1*t1-->1 second 330=(s1+5)(t1-0.5)--->2 solve this two eqn